### v1

# for Scalable MoE Pruning

Abstract
Mixture-of-experts (MoEs) have been adopted for reducing inference costs by sparsely activating experts in Large language models (LLMs). Despite this reduction, the massive number of experts in MoEs still makes them expensive to serve. In this paper, we study how to address this, by pruning MoEs. Among pruning methodologies, unstructured pruning has been known to achieve the highest performance for a given pruning ratio, compared to structured pruning, since the latter imposes constraints on the sparsification structure. This is intuitive, as the solution space of unstructured pruning subsumes that of structured pruning. However, our counterintuitive finding reveals that expert pruning, a form of structured pruning, can actually precede unstructured pruning to outperform unstructured-only pruning. As existing expert pruning, requiring $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, cannot scale for recent MoEs, we propose a scalable alternative with $O(1)$ complexity, yet outperforming the more expensive methods. The key idea is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect. Ours is highly effective– For Snowflake Arctic, a 480B-sized MoE with 128 experts, our method needs only one H100 and two hours to achieve nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to. The code will be made publicly available.
STUN: Structured-Then-Unstructured Pruning
for Scalable MoE Pruning
Jaeseong Lee*, Seung-won Hwangthanks: Work done while visiting Snowflake. Correspond to seungwonh@snu.ac.kr, Aurick Qiao, Daniel Campos, Zhewei Yao, Yuxiong He Snowflake AI Research, Seoul National University*
1 Introduction
Large language models (LLMs) have become the state-of-the-art for various tasks OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Team et al. (2023); Lieber et al. (2024). However, their prohibitive inference cost is becoming a bottleneck to deployment Kaddour et al. (2023), and detrimental to the environment Strubell et al. (2019); Zeng et al. (2023).
Mixture-of-experts (MoE) presents a promising alternative, by sparsely activating a specific subset of parameters, named as experts, to reduce the inference cost. This architecture has been empirically proven effective, in training cost Fedus et al. (2022), and inference cost Du et al. (2022).
Despite the reduction in computational costs, the massive number of parameters remain unchanged, requiring significantly more GPU memory which makes serving large MoE models out of reach for many. To make matters worse, recent MoEs tend to increase the number of experts $n$, resulting in even larger MoEs. For instance, accommodating 56B parameters of Mixtral Jiang et al. (2024) with 8 experts or 132B of DBRX Databricks (2024) with 16 experts, or 480B of Snowflake Arctic Snowflake (2024) with 128 experts, requires an ever-increasing amount of memory and thus more GPUs to host and serve.
Figure 1: GSM8K 5-shot accuracy by pruning Snowflake Arctic, a 480B-sized MoE.
To reduce the number of parameters, unstructured Frantar and Alistarh (2023); Sun et al. (2024), or structured pruning Ma et al. (2023) can be considered. Unstructured pruning allows weight tensors to be sparse anywhere. Structured pruning imposes a restricted pattern on sparsification, such as removing rows, entire weight tensors, or pruning experts in MoE Lu et al. (2024).
Unstructured pruning has been believed to achieve a higher pruning ratio, compared to structured pruning Chen et al. (2022), due to structural restriction. [^1] When comparing Table 1 in Ma et al. (2023) with Table 22 in Sun et al. (2024), the BoolQ performance drops from 71.22 to 61.44, by changing to structured pruning, even allowing less pruning ratio. This seems intuitive, as the solution space of unstructured pruning is a superset of the structured counterpart with structural restrictions. Consequently, an optimal solution that violates the restriction cannot be found from structured pruning, which may result in suboptimal outcomes.
In this paper, we show that Structured-Then-UNstructured pruning (STUN) outperforms unstructured pruning in MoEs. Though it may seem counter-intuitive, we demonstrate that a well-designed structured pruning, which is our expert pruning, would retain the performance after the pruning, and also ensure the pruned network remains robust to unstructured pruning that follows.
First, we argue that MoEs, by design, are ‘robust’ to expert pruning, as MoE training is intended to maintain performance after a number of experts have been pruned. This allows us to design expert pruning to maintain performance similar before and after pruning. To support this argument, we later draw a connection between standard MoE training and targeted dropout Gomez et al. (2019), a training method designed to enhance a network’s robustness to pruning.
Second, we argue that our proposed expert-pruned network remains robust to unstructured pruning. According to Mason-Williams and Dahlqvist (2024), higher kurtosis of weights, or having many outliers in the weight distribution deviating from the normal distribution, suggests more weights can be pruned while maintaining performance, or high robustness of unstructured pruning. We argue that, unlike unstructured pruning, expert (structured) pruning does not decrease kurtosis values, thereby maintaining the network’s robustness to unstructured pruning.
However, existing expert-level pruning for MoE does not scale well over the solution space Lu et al. (2024), requiring an exhaustive combination of experts, leading to $O(\frac{k^{n}}{\sqrt{n}})$ GPU calls, with $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ is pruning ratio. While this was acceptable in an early MoE work with few experts, it does not scale to recent trends in MoEs with large $n$ Bai et al. (2023); Dai et al. (2024); Snowflake (2024), or even infinity He (2024).
Our distinction is drastically reducing the complexity of expert pruning to $O(1)$, without compromising the performance. The main intuition is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect.
Our contributions can be summarized as follows:
  • We propose STUN, the first counterintuitive solution combining structured and unstructured pruning to break the pruning ratio limits of unstructured pruning, to the best of our knowledge.
  • We show that a well-designed expert pruning would retain the performance after the pruning, and the pruned network remains robust to unstructured pruning.
  • To materialize, we design $O(1)$ expert-level pruning method, which outperforms the previously proposed $O(\frac{k^{n}}{\sqrt{n}})$ solution Lu et al. (2024).
  • Our structured pruning, when followed by unstructured pruning, jointly optimizes pruning to achieve a state-of-the-art compression ratio. For Snowflake Arctic, a 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving no loss up to 40% of sparsity, even for generative tasks such as GSM8K, where state-of-the-art unstructured pruning methods fail.
  • Code will be publicly available.
2 Related Work
2.1 LLM Pruning
LLM pruning can be classified into unstructured and structured pruning Behnke and Heafield (2021). Unstructured pruning involves finding mask tensors to sparsify weight tensors. SparseGPT Frantar and Alistarh (2023) uses the Hessian matrix for second-order Taylor approximation, while GBLM-Pruner Das et al. (2024) and Pruner-Zero Dong et al. (2024) leverage gradients to identify mask tensors. However, as these methods demand substantial GPU memory for LLMs, we consider two recent works with higher memory efficiency and strong performance as our baselines: Wanda Sun et al. (2024) evaluates the importance of neurons in each layer by its weight multiplied by the activation value, removing those with low scores. While Wanda assumes a uniform pruning ratio across layers, OWL Yin et al. (2024) probes the optimal pruning ratio per layer, given the pruning budget.
Structured pruning, on the other hand, imposes constraints on the sparsification pattern, such as removing rows, columns, or even entire weight tensors. Early methods that involve pruning attention heads Voita et al. (2019); Shim et al. (2021); Zhang et al. (2021), rows Gong et al. (2022), entire dense layers Liang et al. (2021), or whole transformer blocks Fan et al. (2019); Li et al. (2020) fall under this category. Recent works have applied structured pruning for LLMs Ma et al. (2023); Cheng et al. (2024); Gao et al. (2024); Zhang et al. (2024); Dery et al. (2024), but without fine-tuning, these methods generally underperform compared to unstructured pruning.
Our distinction is to introduce a new class of pruning—structured-then-unstructured pruning—and demonstrate its significant advantages for MoEs, surpassing the performance of unstructured-only pruning. This approach differs from previous methods that combine structured and unstructured pruning Kurtic et al. (2022), which have not succeeded in outperforming unstructured pruning.
2.2 Expert Pruning
Early work on expert pruning was domain-specific, such as in translation MoEs, by keeping most activated experts Kim et al. (2021), or pruning based on gate statistics Koishekenov et al. (2023).
Lu et al. (2024) introduced a domain-agnostic expert pruning, searching for the best combination of experts to reduce the reconstruction loss, and quantify their criticality in output prediction.
Our distinction is eliminating the need for expensive combination enumeration, reducing the GPU calls from $O(\frac{k^{n}}{\sqrt{n}})$ to $O(1)$.
2.3 Pruning Robustness
Robustness in post-hoc pruning is quantified by whether performance is maintained after pruning. Training methods aimed at enhancing this robustness include applying group lasso Wen et al. (2016); Behnke and Heafield (2021), L1 Han et al. (2015); Liu et al. (2017), or L0 Louizos et al. (2018) regularization to the weights, as well as learning a sparse scaling factor Huang and Wang (2018); Li et al. (2020); Gong et al. (2022). Additionally, Targeted Dropout Gomez et al. (2019) demonstrates that applying stochastic dropout to less important parts of the network can increase robustness to pruning.
Meanwhile, kurtosis of weights Mason-Williams and Dahlqvist (2024) has been used as a proxy of robustness, stating networks with higher weight kurtosis can tolerate higher unstructured pruning ratios.
Our contribution lies in connecting these findings for designing expert pruning. We argue that MoE training inherently enhances robustness to expert-level pruning, and this coarse structured pruning subsequently makes the remaining network more resilient to unstructured pruning. In section 5, we argue the former, based on the resemblance of MoE training and targeted dropout, and the latter, inspired by Mason-Williams and Dahlqvist (2024).
3 Preliminaries: MoE
As a promising alternative to large language models, which incur prohibitive inference costs, MoE employs a multitude of specialized experts. In each forward pass, MoE selectively activates specific experts conditioned on input tokens, thereby reducing the train and inference costs.
We now formally describe the behavior of an MoE. An MoE layer $M$ consists of experts $E_{i}=E(x;\theta_{i})$, where $\theta_{i}$ represents the parameters of expert $E_{i}$, and a router layer $r$. Each expert $E$ typically follows the same MLP architecture.
First, the router layer selects which experts to sparsely activate based on the current input token, and provides the coefficients $r(x)\in\mathbb{R}^{n}$ for linear combination of selected expert outputs. The coefficients $r(x)$ and the top-k indices of experts $\mathcal{T}$ are formulated as follows:
$r(x)=\mathrm{softmax}(Wx)$ (1)
$\mathcal{T}=\mathrm{topk}(r(x))$ (2)
where $W$ is the learnable weight matrix for router $r$.
Next, these coefficients are used for the linear combination of expert outputs:
$M(x;\mathbf{\theta})=\sum_{i\in\mathcal{T}}r_{i}(x)E(x;\theta_{i})$ (3)
4 Structured-Then-UNstructured Pruning (STUN)
4.1 Overview
Essentially, STUN first performs structured (expert) pruning, until the loss is negligible, then followed by unstructured pruning, such as OWL Yin et al. (2024) or Wanda Sun et al. (2024).
Our key contribution is to replace combinatorial loss with $O(1)$ expert pruning method, by leveraging latent cluster structure among experts, based on behavioral similarity.
Specifically, we find clusters of similar experts layer by layer, yielding a total of $\phi nl$ clusters in the whole MoE, where $\phi$ is the pruning ratio, $n$ is the number of experts in each layer, and $l$ is the number of layers in MoE. Then we greedily prune every expert but one representative per each cluster.
Later sections show why our greedy pruning is as effective as its combinatorial counterpart, and how it effectively connects to MoE training beforehand, and unstructured pruning afterwards.
4.2 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
Now we will derive our algorithm in section 4.1 formally, starting from the conventional goal of pruning– minimizing the reconstruction loss. Reconstruction loss has been employed to assess how closely the pruned model $\theta-\theta_{S}$ without expert set $S$ mirrors the behavior of the unpruned $\theta$  Lu et al. (2024). Formally, this loss is quantified by the Frobenius norm of the difference between the original output $M(x;\theta)$ and the output of pruned layer $M(x;\theta-\theta_{S})$, denoted as $\mathcal{E}_{S}$.
$\mathcal{E}_{S}=\|M(x;\theta)-M(x;\theta-\theta_{S})\|_{F}$ (4)
where $x$ is the whole input we consider. The objective of pruning is to explore all possible combinations of experts, (${n}\choose{|S|}$, to determine the expert set $S$ that minimizes $\mathcal{E}_{S}$.
While such an exhaustive search is feasible for smaller models like Mixtral Jiang et al. (2024), which contains only 8 experts, it becomes prohibitive for recent MoEs featuring a massive number of experts.
To elaborate, deciding which experts to prune using Eq. 4 for $|S|=\phi n$ requires (${n\choose{|S|}}\approx O(\frac{k^{n}}{\sqrt{n}})$ forward passes according to Stirling’s approximation, where $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ represents the pruning ratio. [^2] In detail, 23951146041928082866135587776380551750 forward passes per layer at minimum, for massive MoEs with $n=128$ experts such as Snowflake Arctic Snowflake (2024). Our distinction is to lower the computation to $O(1)$, without compromising the performance– In fact, as we will elaborate later, we outperform the combinatorial objective.
4.3 Towards $O(n)$: Probabilistic Interpretation
As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ as:
$\textrm{argmax}_{S}P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ (5)
where $s_{i}$s are the experts included in the expert set $S$, and $P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ is the joint probability of pruning $S$.
Section 4.2 corresponds to enumerating joint probability from all combinations, requiring ($n\choose{|S|}$ different values, which is compute-intensive. When chain rule is applied, Eq. 5 can be reformulated as follows:
$\textrm{argmax}_{S}\prod_{k}P(X_{k}=s_{k}|X_{1}=s_{1},\\ \cdots,X_{k-1}=s_{k-1})$ (6)
We propose a greedy optimization for Eq. 6– We decompose the multiplication of Eq. 6 and at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability. For simplicity, we will omit $X_{k}$s from this point on.
As our goal is finding the argmax of the probabilities as in Eq. 6, estimating the rank between them is sufficient, rather than evaluating exact values. Such rank estimation can benefit from the latent structure among experts, specifically, a cluster of similar experts in MoE, enabling $P(X_{k}|s_{1},\cdots,s_{k-1})$ calculation without chain-rule multiplications in Eq. 6.
Assume we know oracle clusters, $c(E_{i})$, where $c$ is the mapping from an expert to a set of similarly behaving experts identified from the latent clusters. When we have knowledge of similar experts, for example, $c(E_{i})=c(E_{j})=\{E_{i},E_{j}\}$ indicating $E_{i}$ and $E_{j}$ are highly similar, we will decide not to prune $E_{i}$ if $E_{j}$ is already pruned. That is, if $c(E_{i})\subseteq S_{k}$ then $P(X_{k}=E_{i}|S_{k-1})$ should be lowered by some value $p$, to guide the model against pruning. Moreover, $P(E_{i}|S_{k-1})$ should be larger, or rank higher, otherwise.
To generalize, we will cluster similar experts. Once the cluster of similar experts is finalized, we assign the value $P(E_{i}|S_{k-1})$, as follows:
$P(E_{i}|S_{k-1})=\begin{cases}P(E_{i})-p&c(E_{i})\subseteq S_{k}\\ P(E_{i})&\mathrm{otherwise}\end{cases}$ (7)
We set $p$ as a constant for simplicity. This enables the calculation of all $P(E_{i}|S_{k-1})$ in Eq. 6 from $P(E_{i})$s, which needs only $n$ forwards in total.
Clustering the Similar Experts
Our remaining task is to obtain cluster information $c$: One signal is pairwise behavioral similarity $b_{i,j}$, from the pretrained weights $W$ at a minimal cost. Suppose two rows $W_{i}\approx W_{j}$ are similar; then $r_{i}(x)\approx r_{j}(x)$, meaning $E_{i},E_{j}$ tend to be selected by similar inputs, implying similar expertise. Thus, the behavioral similarity $b_{i,j}$ between two experts $E_{i},E_{j}$ can be obtained as follows:
$b_{i,j}=-\|W_{i}-W_{j}\|_{F}$ (8)
Next, we generalize pairwise similarity into clusters of experts, such that experts in each cluster $C_{l}$ are highly similar to its representative $\mu_{l}$. Formally, the objective of clustering is to minimize the sum of squared errors between $\mu_{l}$ and experts $E_{i}$ in the cluster:
$\sum_{i\in C_{l}}\sum_{l}\|W_{i}-\mu_{l}\|^{2}$ (9)
which is an NP-hard problem Megiddo and Supowit (1984); Dasgupta (2008); Aloise et al. (2009).
Practically, we found that the agglomerative clustering algorithm Sneath and Sokal (1973) performs well. [^3] We tried other clustering algorithms in the Appendix. Specifically, clusters are initialized as individual experts and then iteratively merged, with a termination condition that prevents the experts within each cluster from being too dissimilar. This condition is tuned based on the desired pruning ratio.
Lastly, if we allow inference on some data, we can improve Eq. 8 with coactivation statistics $a_{i,j}$, which measure the frequency with which $E_{i},E_{j}$ are selected simultaneously. [^4] We normalize $a_{i,j}$ by dividing it with the total coactivations in one layer. However, these coactivation statistics depend on the given data, whose distribution may differ from the test data. Therefore we combine the two as follows:
$b_{i,j}=-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$ (10)
We recap the algorithm in the Appendix (Alg 1).
4.4 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
1st-order Taylor Approximation
While the $O(n)$ approach immensely reduces the cost to obtain the probability distribution, we can further reduce the number of forward passes needed by precalculating $P(E_{i})$s in Eq. 7.
The key idea is approximating $E_{i}$’s reconstruction loss value $\mathcal{E}_{i}=\|M(x;\theta)-M(x;\theta-\theta_{i})\|_{F}$, and assigning $P(E_{i})$ as some high value $L$ if the reconstruction loss $\mathcal{E}_{i}$ is lowest. This neatly estimates the rank between $P(E_{i})$s.
Though $\mathcal{E}_{i}$ can be approximated via conventional 2nd-order reconstruction methods Hassibi and Stork (1992); Frantar and Alistarh (2023), the size of the hessian matrix increases quadratically with the number of experts, which often yields out-of-memory errors.
To address this, we propose using a 1st order Taylor approximation. To rank the reconstruction loss values, we consider approximating the reconstruction loss when replacing the output from $\theta_{i}$ with some specific expert $\theta_{C}$ in $C=c(E_{i})$ as follows:
$\mathcal{E}_{i}=\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ (11)
As the convention of 2nd order Taylor approximation Hassibi and Stork (1992); Frantar and Alistarh (2023), we assume the parameters are near a local minimum. Thus, with a small constant $\gamma$, $\|E^{\prime}(\theta_{i})\|<\gamma$, leading to:
$\sum_{i}\mathcal{E}_{i}<\sum_{i}\gamma\|\theta_{i}-\theta_{C}\|^{2}$ (12)
whose upper bound in the right-hand side can be minimized when $\theta_{C}=\overline{\theta_{i}}$, where $\bar{~{}}$ denotes the average.
Therefore, the expert closest to $\bar{\theta_{i}}$ within each cluster has the highest priority to be retained. We assign $\mathcal{E}_{i}$ a large number $L>p$ if $E_{i}$ is the closest to $\bar{\theta_{i}}$ from the corresponding cluster $c(E_{i})$, and set it to zero otherwise. The same greedy algorithm is applied to optimize Eq. 6.
Selective Reconstruction of Experts
While letting $\theta_{C}$ as the expert closest to $\bar{\theta_{i}}$ successfully minimizes $\sum_{i}\mathcal{E}_{i}$, sometimes we can minimize them further, by replacing the weight of the closest expert $\theta_{C}$ to $\bar{\theta_{i}}$. However, blindly doing so is suboptimal, as there is another kind of error to consider. The decision boundaries of the next layer are accustomed to the output of $\{E(x;\theta_{i})\}_{i=1}^{|C|}$, but changing the output as $E(x;\theta_{C})=E(x;\bar{\theta_{i}})$ could result in a distribution that the model is unfamiliar with. This potential error, which we denote as $\mathcal{E}_{d}$, would be minimized if $\theta_{C}\in\{\theta_{i}\}_{i=1}^{|C|}$.
To balance these two types of errors, we selectively decide whether to reconstruct. We observe that $\sum_{i}\mathcal{E}_{i}$ increases if the total number of clusters in a layer decreases, as this would introduce more $\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ terms. Therefore, if the total number of clusters is below a threshold $\kappa$, we use $\theta_{C}=\overline{\theta_{i}}$ to minimize $\sum_{i}\mathcal{E}_{i}$. Otherwise, we set $\theta_{C}$ as the expert within the cluster $\{\theta_{i}\}_{i=1}^{|C|}$ closest to the $\overline{\theta_{i}}$, to minimize $\mathcal{E}_{d}$. The router weight reconstruction is done similarly, following its corresponding expert. The final algorithm is summarized in the Appendix (Alg 2).
5 Robustness of Structured-Then-Unstructured Pruning
In this section, we deliver the key insights into why STUN works: MoEs are inherently robust to expert pruning, and the expert-pruned MoE is robust to subsequent post-hoc unstructured pruning.
MoE Robustness to Structured (Expert) Pruning
To support this, we find a resemblance between the MoE (Eq. 3) and the targeted dropout Gomez et al. (2019). Targeted dropout achieves a higher post-hoc pruning ratio by stochastically dropping out unimportant units or weights during training. Formally, in each training step, $M(x;\theta)$ is modified as follows:
$M(x;\theta-U(\theta)\odot D)$ (13)
where $U(\theta)$ is a function that extracts only an unimportant part of a given weight $\theta$, and $D$ is a stochastic mask to ensure $U(\theta)$ also has a chance to be updated. Gomez et al. (2019) designed $U(\theta)$ as the bottom $|\theta|-k$ elements sorted by their absolute values, and $D\sim Bernoulli(\alpha)$ where $\alpha$ is the dropout ratio. After training, the performance is better retained when $U(\theta)$ is pruned.
We recognize the resemblance between the MoE (Eq. 3) and the targeted dropout (Eq. 13). If we interpret $U(\theta)$ as $\bigcup_{i\notin\mathcal{T}}E(\theta_{i})$, then $M(x;\theta-U(\theta))$ resembles Eq. 3. Moreover, just as $D$ is added in Eq. 13 to ensure that $U(\theta)$ is updated, MoE ensures that $U(\theta)$ is updated through the stochastic nature of $r(x)$.
Therefore, after training, MoE is expected to be robust to pruning $U(\theta)$, which, in this context, corresponds to pruning some experts $E(\theta_{i})$s. However, identifying $U(\theta)$ is not as straightforward as in Gomez et al. (2019). In MoE training, $U(\theta)$ is conditioned on each token, such that we need to marginalize token dependency, by identifying unimportant experts across a broad range of inputs. This challenge is connected to the problem of finding $S$ in Eq. 4, which minimizes the reconstruction loss.
Expert-Pruned MoE Robustness to Unstructured Pruning
Robustness to unstructured pruning can be estimated by the kurtosis of weights Mason-Williams and Dahlqvist (2024). Kurtosis is expressed as follows:
$K(\theta)=E\left[\left(\frac{\theta-\mu}{\sigma}\right)^{4}\right]$ (14)
Suppose the weight of experts $\theta$ follow a zero-meaned Gaussian distribution $\mathcal{N}$. Unstructured pruning Sun et al. (2024); Yin et al. (2024); Das et al. (2024); Dong et al. (2024), which tends to remove near-zero weights, [^5] The importance score of unstructured pruning typically increases as the absolute value of the weight increases. would shift the distribution closer to a bimodal symmetric distribution, whose kurtosis is minimum Darlington (1970). As a result, unstructured pruning would lower the kurtosis value, leaving less margin for further unstructured pruning.
In contrast, coarse structured pruning, such as expert pruning, is less likely to decrease the kurtosis value, since the assumption $\theta\sim\mathcal{N}$ still holds for remaining experts. This implies that expert pruning preserves the robustness of unstructured pruning, unlike applying unstructured pruning with a similar pruning ratio.
6 Experiments
model sparsity method GSM8K Avg(→) ARC-c ARC-e HellaSwag MMLU
Arctic 0% unpruned 70.74 68.33 56.91 84.60 66.94 64.86
40% STUN (w/ OWL) 70.28 67.66 57.68 83.29 64.94 64.75
OWL 63.76 67.35 56.74 84.13 65.08 63.47
STUN (w/ Wanda) 69.60 67.64 57.25 83.63 64.86 64.81
Wanda 64.59 67.54 57.00 84.64 65.19 63.32
65% STUN (w/ OWL) 43.97 62.67 51.54 80.01 59.91 59.24
OWL 13.42 56.68 44.37 76.64 53.69 52.02
Mixtral-8x7B (Instruct) 65% STUN (w/ OWL) 25.09 60.34 48.12 78.79 54.05 60.39
OWL 1.29 45.20 24.15 49.79 49.27 57.60
Mixtral-8x22B (Instruct) 70% STUN (w/ OWL) 30.78 60.20 47.95 77.86 55.41 59.56
OWL 19.64 57.74 45.48 76.60 52.47 56.42
Table 1: Comparison between STUN and the baselines across various models.
model sparsity method cost ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
Mixtral-8x7B (Instruct) 0% unpruned 62.20 87.04 88.50 67.59 68.87 36.60 72.20 76.87 69.98
25% Ours $O(1)$ 59.30 85.44 88.13 64.42 64.52 35.40 71.84 75.37 68.05
Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 58.19 84.89 87.34 65.24 62.47 35.60 70.04 75.85 67.45
Mixtral-8x7B 0% unpruned 57.17 84.01 85.35 64.88 67.88 35.00 70.40 75.93 67.58
25% Ours $O(1)$ 52.73 81.82 83.09 60.84 63.34 31.60 68.59 72.69 64.34
Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 51.62 81.94 83.64 61.60 58.72 33.00 67.87 75.37 64.22
Table 2: Comparing the first component of STUN, the proposed expert pruning, with other baselines.
(a) Arctic (128x3.66B)
(b) Mixtral-8x7B
(c) Mixtral-8x22B
Figure 2: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various MoEs.
In this section, we aim to address the following research questions:
  • RQ1: Does STUN outperform unstructured pruning alone?
  • RQ2: Does the expert pruning we propose outperform existing baselines?
  • RQ3: Does STUN favor recent MoE trends with a large number of small experts?
  • RQ4: Is each component of STUN essential for achieving performance gains?
  • RQ5: Does STUN generalize to non-MoE models as well?
6.1 Experimental Settings
We use Snowflake Arctic Snowflake (2024) as a representative large MoE, with a total of 480B parameters and 128 experts. To compare our method with previous works Lu et al. (2024), we also experiment with Mixtral Jiang et al. (2024).
Tasks and Datasets
In contrast to previous unstructured pruning studies Sun et al. (2024); Yin et al. (2024), we also evaluate on the NLG task, GSM8K Cobbe et al. (2021), where maintaining performance proves to be much more challenging (see Table 1). We further assess performance on four NLU tasks– ARC-challenge and ARC-easy Clark et al. (2018), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2021). When comparing with expert pruning methods, following previous work Lu et al. (2024), we also conduct a zero-shot evaluation on BoolQ Wang et al. (2019), OpenBookQA Mihaylov et al. (2018), RTE Wang et al. (2018), WinoGrande Sakaguchi et al. (2021).
To provide some data for inference, we employ the C4 dataset Raffel et al. (2020), following the baselines Yin et al. (2024); Sun et al. (2024); Lu et al. (2024).
Implementation Details
We explore the values of $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for Snowflake Arctic, the largest MoE in our experiments, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, meaning no GPU calls are required for expert pruning. We evaluate methods using the lm-evaluation-harness Gao et al. (2021). Due to the model size, we use 4-bit quantization Dettmers et al. (2023) for experiments with Mixtral-8x22B and Arctic. The expert pruning ratios are set to 20%, 12.5%, and 10% for Snowflake Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively. We use $\kappa=3$ for selective reconstruction. More detailed implementation information is provided in the Appendix.
6.2 Experimental Results
6.2.1 RQ1: STUN Outperforms Unstructured Pruning
Table 1 describes that our proposed (STUN) significantly outperforms the unstructured pruning methods. We emphasize that we use the same unstructured pruning approach for all for fair comparison.
For example, with 40% of sparsity for the Arctic, STUN neatly retains the original GSM8K performance, while unstructured pruning results in a noticeable performance drop. This is consistent for different unstructured pruning methods, Wanda, as well. For 65% of sparsity for Mixtral-8x7B-Instruct, STUN’s GSM8K performance is nearly 20 times better than that of unstructured pruning. In the ARC-challenge, the unstructured pruning performance falls below the random-guess accuracy of 25.02 Clark et al. (2018), whereas STUN maintains a significantly higher performance, achieving twice the score.
6.2.2 RQ2: Our $O(1)$ Expert Pruning Outperforms Existing $O(\frac{k^{n}}{\sqrt{n}})$
Table 2 shows that our proposed $O(1)$ expert pruning method is highly effective, outperforming the previous $O(\frac{k^{n}}{\sqrt{n}})$ solution. This is because we derive the latent structure from the pretrained MoE, while the previous work Lu et al. (2024) relies solely on the given calibration data. This validates our design of $O(1)$ in section 4.
6.2.3 RQ3: STUN Favors Large Number of Small Experts
Figure 2 illustrates the trend of STUN in 3 different MoEs. The performance gap between STUN and unstructured pruning increases as the MoE has more experts with small sizes (from (c) to (a)). This is because having more experts, rather than having fewer but larger ones, provides greater flexibility to our expert pruning. Notably, MoEs with a large number of small experts are favored in recent models He (2024).
Cluster (§$\S$ 4.3) Reconstruct (§$\S$ 4.4) LM-eval Avg
Ours Ours 59.58
DSatur Ours 58.59
Ours Always 57.60
Ours Never 59.22
Table 3: Ablation experiments for the first component of STUN, the proposed expert pruning.
6.2.4 RQ4: Ablation
To validate our design of expert pruning in sections 4.3 and 4.4, we evaluate alternative approaches to expert-prune Mixtral-8x7B at 50% sparsity. Table 3 confirms that our design choices are valid. Our agglomerative clustering algorithm outperforms the DSatur algorithm, an alternative clustering algorithm we discuss in the Appendix. Additionally, selective reconstruction proves superior to always or never reconstructing, as shown in the last two rows. Detailed evaluation results are provided in the Appendix.
6.2.5 RQ5: STUN Outperforms Unstructured Pruning in non-MoEs
To investigate whether STUN is generalizable to non-MoE as well, we employ a state-of-the-art structured pruning algorithm for non-MoE model, namely, LLM-surgeon van der Ouderaa et al. (2024) with 5% sparsity before performing unstructured pruning, which is OWL in our case. Figure 3 illustrates that such STUN outperforms unstructured pruning.
(a) Llama-2 7B
(b) Llama-2 13B
Figure 3: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various non-MoEs.
7 Conclusion
In this paper, we proposed STUN for MoEs, which outperforms unstructured-only pruning. Although this may seem counterintuitive, we provide both theoretical and empirical evidence demonstrating why designing expert pruning before unstructured pruning is beneficial. Since existing expert pruning requires $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, we proposed $O(1)$ complexity, which outperforms the more expensive approaches. Our experimental results show STUN  is highly effective– For 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to.
Limitation
Since our method utilizes unstructured pruning in the second stage, we share the same disadvantages with unstructured pruning, that is, it may need specialized hardware for acceleration. However, it is shown that general-purpose hardware, such as CPU, can successfully accelerate unstructure-pruned networks NeuralMagic (2021). Therefore we have the same potential as existing unstructured pruning methods– with the support from the community, our method will scale up to existing accelerators.
Appendix
Algorithm of $O(1)$ Expert Pruning
Algorithms 1 and 2 describe the expert pruning we proposed. Note that we do not introduce any GPU inference in both two algorithms.
Algorithm 1 Expert Clustering Algorithm
$l\leftarrow$ Number of layers
$n\leftarrow$ Number of experts per layer
$\{a_{i,j}\}_{i,j}\leftarrow$ Coactivation statistics of $E_{i},E_{j}$ for every layer
$\lambda_{1},\lambda_{2}\leftarrow$ Hyperparameter for behavioral similarity
$t\leftarrow$ Threshold to determine pruning ratio
[…]
     $W\leftarrow$ Router weight of layer $m$
     $\{a_{i,j}\}_{i,j}\leftarrow$ Coactivation statistics of layer $m$
[…]
              $b_{i,j}\leftarrow-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$
         end for
     end for
[…]
         $c(E_{i})\leftarrow\{E_{i}\}$
     end for
     while $\min_{i,j}b_{i,j}<t$ do
         ${d,e}\leftarrow\mathrm{argmin}_{i,j}b_{i,j}$
         $m_{d}\leftarrow\max_{i\in c(E_{e})}b_{d,i}$
         $m_{e}\leftarrow\max_{i\in c(E_{d})}b_{i,e}$
         if $c(E_{d})\neq c(E_{e})\wedge\max(m_{d},m_{e})<t$ then
              $c(E_{d})=c(E_{e})\leftarrow c(E_{d})\cup c(E_{e})$
         end if
         $b_{d,e}\leftarrow\infty$triangleright Mark as visited
     end while
end for
return $c$
Algorithm 2 Our $O(1)$ Expert Pruning
$l\leftarrow$ Number of layers
$n\leftarrow$ Number of experts per layer
$c\leftarrow$ The mapping from expert to cluster of the similar experts
$\kappa\leftarrow$ Threshold for selective reconstruction
[…]
     $r(m)=[~{}]$
     $A\leftarrow\{c(E_{1}),\cdots,c(E_{n})\}$
     for $C$ in $A$ do
         $\bar{\theta_{i}}\leftarrow\frac{1}{|C|}\sum_{i\in C}\theta_{i}$
         if $|A|<\kappa$ then
              $\theta_{C}\leftarrow\bar{\theta_{i}}$triangleright Reconstruct
         else
              $\theta_{C}\leftarrow\min_{\theta_{j}\in C}\|\theta_{j}-\bar{\theta_{i}}\|_{F}$
         end if
         $r(m).append(\theta_{C})$
     end for
end for
return $r$
Implementation Details
We probe $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for the Snowflake Arctic, which is the biggest MoE we deal with, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, which means no GPU calls is needed for expert pruning. To get coactivation values $a_{i,j}$, we utilize 1000 samples from the C4 dataset, each of which has 2048 sequence length. We evaluate on lm-evaluation-harness Gao et al. (2021) We use 4bit quantization Dettmers et al. (2023) for experiments with Mixtral-8x22B and Arctic, due to their model size. We use 20%, 12.5%, and 10% for Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively, as the expert pruning ratio for STUN. These are the maximum values among 10, 12.5, 20, 25, and 35%, with minimum performance loss. We use $\kappa=3$ for selective reconstruction. For Wanda and OWL, we use 128 C4 samples following the original papers Yin et al. (2024); Sun et al. (2024), while we use 4096 for sequence length. For OWL, we use the default setting, $M=5,\lambda=0.08$.
All experiments are conducted on H100 80GB GPUs, with a maximum of 4. Each evaluation is done within 4 hours, and each unstructured pruning requires less than 2 hours on one GPU. Evaluation is done only once, since we introduce no randomness in our experiment.
Other Clustering Algorithms
We also considered DSatur Brélaz (1979) as a clustering algorithm for Eq. 12, converting into clique-partitioning in a graph where edge $e_{i,j}$ connected if two experts are similar enough as follows,
$e_{i,j}=\begin{cases}1&b_{i,j}>=t_{DSatur}\\ \infty&\mathrm{otherwise}\end{cases}$ (15)
where $t_{DSatur}$ is some threshold to control the pruning ratio of MoE.
Ablation Studies Results
Detailed per-task performance of ablation studies are described in Tables 4, 5.
ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
DSatur 45.99 75.38 80.76 54.62 45.06 29.00 66.79 71.11 58.59
Ours 45.73 75.13 83.46 54.55 53.29 31.20 62.45 70.80 59.58
Table 4: Our agglomerative clustering algorithm is better than the alternative.
ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
Always reconstruct ($\kappa=8$) 42.92 74.20 82.42 54.66 50.09 28.40 55.96 72.14 57.60
No reconstruct ($\kappa=0$) 45.22 75.21 82.45 54.53 52.31 30.00 62.82 71.19 59.22
Ours ($\kappa=3$) 45.73 75.13 83.46 54.55 53.29 31.20 62.45 70.80 59.58
Table 5: Selective reconstruction outperforms the baselines.

### v2

@@
 Abstract
-Mixture-of-experts (MoEs) have been adopted for reducing inference costs by sparsely activating experts in Large language models (LLMs).
-Despite this reduction, the massive number of experts in MoEs still makes them expensive to serve.
-In this paper, we study how to address this, by pruning MoEs.
-Among pruning methodologies, unstructured pruning has been known to achieve the highest performance for a given pruning ratio, compared to structured pruning, since the latter imposes constraints on the sparsification structure.
-This is intuitive, as the solution space of unstructured pruning subsumes that of structured pruning.
-However, our counterintuitive finding reveals that expert pruning, a form of structured pruning, can actually precede unstructured pruning to outperform unstructured-only pruning.
-As existing expert pruning, requiring $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, cannot scale for recent MoEs, we propose a scalable alternative with $O(1)$ complexity, yet outperforming the more expensive methods.
-The key idea is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect.
-Ours is highly effective– For Snowflake Arctic, a 480B-sized MoE with 128 experts, our method needs only one H100 and two hours to achieve nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to.
-The code will be made publicly available.
+Mixture-of-experts (MoEs) have been adopted to reduce inference costs by sparsely activating experts in large language models (LLMs).
+Despite these reductions, the massive number of parameters in MoEs still makes them expensive to serve.
+Conventionally, unstructured or structured pruning has been considered to reduce the number of parameters.
+Our key contribution is exploring the interpolation between structured and unstructured pruning, to propose a novel structured-then-unstructured (STUN) approach outperforming both structured and unstructured pruning, especially for MoEs.
+In the first stage, we show a scalable expert pruning with O(1) forward pass, unlike existing work requiring $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts that cannot scale for recent MoEs with hundreds of experts.
+We then show our expert-pruned MoEs are robust to unstructured pruning to follow.
+Experiments on Snowflake Arctic and Mixtral show that our proposal is highly effective– For Snowflake Arctic, a 480B-sized MoE with 128 experts, our method needs only one H100 and two hours to achieve nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art structured or unstructured pruning methods fail.
+The code is publicly available.
+[^1] https://github.com/thnkinbtfly/STUN
 STUN: Structured-Then-Unstructured Pruning
@@
 1 Introduction
-Large language models (LLMs) have become the state-of-the-art for various tasks OpenAI (2023); Touvron et al.
+Large language models (LLMs) have become state-of-the-art for various tasks OpenAI (2023); Touvron et al.
 (2023); Jiang et al.
-(2023); Team et al.
 (2023); Lieber et al.
@@
 (2022).
-Despite the reduction in computational costs, the massive number of parameters remain unchanged, requiring significantly more GPU memory which makes serving large MoE models out of reach for many.
-To make matters worse, recent MoEs tend to increase the number of experts $n$, resulting in even larger MoEs.
+Despite these reductions, the massive number of parameters remains unchanged, requiring significantly more GPU memory, which makes serving large MoE models challenging for many.
+Additionally, recent MoEs tend to increase the number of experts $n$, resulting in even larger MoEs.
 For instance, accommodating 56B parameters of Mixtral Jiang et al.
-(2024) with 8 experts or 132B of DBRX Databricks (2024) with 16 experts, or 480B of Snowflake Arctic Snowflake (2024) with 128 experts, requires an ever-increasing amount of memory and thus more GPUs to host and serve.
-Figure 1: GSM8K 5-shot accuracy by pruning Snowflake Arctic, a 480B-sized MoE.
+(2024) with 8 experts or 132B of DBRX Databricks (2024) with 16 experts, or 480B of Snowflake Arctic Snowflake (2024) with 128 experts, requires an ever-growing amount of memory and more GPUs to serve.
+Figure 1: GSM8K 5-shot accuracy by pruning Mixtral-8x7B-Instruct by 50% of sparsity. We probe interpolation of structured and unstructured pruning, by varying the ratio of structured pruning.
 To reduce the number of parameters, unstructured Frantar and Alistarh (2023); Sun et al.
@@
 (2023) can be considered.
-Unstructured pruning allows weight tensors to be sparse anywhere.
-Structured pruning imposes a restricted pattern on sparsification, such as removing rows, entire weight tensors, or pruning experts in MoE Lu et al.
-(2024).
-Unstructured pruning has been believed to achieve a higher pruning ratio, compared to structured pruning Chen et al.
-(2022), due to structural restriction.
-[^1] When comparing Table 1 in Ma et al.
-(2023) with Table 22 in Sun et al.
-(2024), the BoolQ performance drops from 71.22 to 61.44, by changing to structured pruning, even allowing less pruning ratio.
-This seems intuitive, as the solution space of unstructured pruning is a superset of the structured counterpart with structural restrictions.
-Consequently, an optimal solution that violates the restriction cannot be found from structured pruning, which may result in suboptimal outcomes.
-In this paper, we show that Structured-Then-UNstructured pruning (STUN) outperforms unstructured pruning in MoEs.
-Though it may seem counter-intuitive, we demonstrate that a well-designed structured pruning, which is our expert pruning, would retain the performance after the pruning, and also ensure the pruned network remains robust to unstructured pruning that follows.
-First, we argue that MoEs, by design, are ‘robust’ to expert pruning, as MoE training is intended to maintain performance after a number of experts have been pruned.
-This allows us to design expert pruning to maintain performance similar before and after pruning.
-To support this argument, we later draw a connection between standard MoE training and targeted dropout Gomez et al.
-(2019), a training method designed to enhance a network’s robustness to pruning.
-Second, we argue that our proposed expert-pruned network remains robust to unstructured pruning.
-According to Mason-Williams and Dahlqvist (2024), higher kurtosis of weights, or having many outliers in the weight distribution deviating from the normal distribution, suggests more weights can be pruned while maintaining performance, or high robustness of unstructured pruning.
-We argue that, unlike unstructured pruning, expert (structured) pruning does not decrease kurtosis values, thereby maintaining the network’s robustness to unstructured pruning.
-However, existing expert-level pruning for MoE does not scale well over the solution space Lu et al.
-(2024), requiring an exhaustive combination of experts, leading to $O(\frac{k^{n}}{\sqrt{n}})$ GPU calls, with $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ is pruning ratio.
+Unstructured pruning allows weight tensors to be sparse anywhere, while structured pruning imposes patterns on sparsification, such as removing rows, entire weight tensors Ma et al.
+(2023), or pruning experts in MoE Lu et al.
+(2024a).
+In this paper, we propose the interpolation of two, Structured-Then-UNstructured pruning (STUN).
+Figure 1 motivates our interpolated method, where unstructured- or structured-only, $x=0$ and $x=1$, respectively, is outperformed by the peak, combining both.
+Our first phase, expert pruning, by leveraging model-inherent expert for pruning, significantly outperforms row/column-level structured pruning (Figure 1 blue; Ma et al., 2023) However, existing expert-level pruning for MoE (Figure 1 grey; Lu et al., 2024a) often does not scale well over the solution space, requiring an exhaustive combination of experts, leading to $O(\frac{k^{n}}{\sqrt{n}})$ GPU calls, with $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ is sparsity Lu et al.
+(2024a).
 While this was acceptable in an early MoE work with few experts, it does not scale to recent trends in MoEs with large $n$ Bai et al.
@@
 (2024); Snowflake (2024), or even infinity He (2024).
-Our distinction is drastically reducing the complexity of expert pruning to $O(1)$, without compromising the performance.
+Our distinction is drastically reducing the number of GPU calls to $O(1)$, without compromising the performance.
 The main intuition is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect.
+The second contribution is allowing unstructured phase to follow, to consider both inter- and intra-expert sparsity (Table 1).
+STUN removes redundant experts by expert-level structured pruning first, then desires fine-grained sparsity within individual experts.
+[^2] From now on, we will define sparsity as the number of pruned parameters divided by the total number of parameters in the original model.
+We support STUN with the findings of Mason-Williams and Dahlqvist (2024), which show that higher kurtosis in the weight distribution (indicating many outliers) suggests more weights can be pruned while maintaining performance, highlighting the robustness of unstructured pruning.
+We argue that expert-level pruning does not reduce kurtosis, thereby preserving the network’s resilience to unstructured pruning.
+Inter-Expert Sparsity Intra-Expert Sparsity Figure 1
+Lu et al., 2024a ($\bullet$)
+LLM-Pruner ($\bullet$), Wanda, OWL ($\bullet$)
+Ours ($\bullet$)
+Table 1: Comparison of existing pruning methods for MoEs.
 Our contributions can be summarized as follows:
-  • We propose STUN, the first counterintuitive solution combining structured and unstructured pruning to break the pruning ratio limits of unstructured pruning, to the best of our knowledge.
-  • We show that a well-designed expert pruning would retain the performance after the pruning, and the pruned network remains robust to unstructured pruning.
-  • To materialize, we design $O(1)$ expert-level pruning method, which outperforms the previously proposed $O(\frac{k^{n}}{\sqrt{n}})$ solution Lu et al. (2024).
-  • Our structured pruning, when followed by unstructured pruning, jointly optimizes pruning to achieve a state-of-the-art compression ratio.
-For Snowflake Arctic, a 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving no loss up to 40% of sparsity, even for generative tasks such as GSM8K, where state-of-the-art unstructured pruning methods fail.
-  • Code will be publicly available.
+  • We propose STUN, the first method to combine structured and unstructured pruning, outperforming both approaches.
+  • Scalable first phase: We design an expert-level pruning method with $O(1)$ GPU calls, outperforming the $O(\frac{k^{n}}{\sqrt{n}})$ solution Lu et al. (2024a).
+  • Justifying the second phase: We show the expert-pruned network remains robust to unstructured pruning to follow.
+  • State-of-the-art efficiency and compression: For Snowflake Arctic (480B, 128 experts), it requires just 1 H100 and two hours, with no backpropagation or fine-tuning needed.
+Compression reaches up to 40% sparsity without compromising performance, even in generative tasks like GSM8K, where unstructured pruning fails.
+We report consistent results for Mixtral models.
 2 Related Work
@@
 Unstructured pruning involves finding mask tensors to sparsify weight tensors.
+Such masking leads to practical speedups in hardware such as CPU NeuralMagic (2021), and ongoing research is actively developing methods to achieve similar speedups on GPUs Mishra et al.
+(2021); Zhao et al.
+(2024).
 SparseGPT Frantar and Alistarh (2023) uses the Hessian matrix for second-order Taylor approximation, while GBLM-Pruner Das et al.
@@
 (2024) leverage gradients to identify mask tensors.
-However, as these methods demand substantial GPU memory for LLMs, we consider two recent works with higher memory efficiency and strong performance as our baselines: Wanda Sun et al.
+However, as these methods demand substantial GPU memory for LLMs, we focus on more memory-efficient approaches, using two recent baselines: Wanda Sun et al.
 (2024) evaluates the importance of neurons in each layer by its weight multiplied by the activation value, removing those with low scores.
-While Wanda assumes a uniform pruning ratio across layers, OWL Yin et al.
-(2024) probes the optimal pruning ratio per layer, given the pruning budget.
+While Wanda assumes a uniform sparsity across layers, OWL Yin et al.
+(2024) probes the optimal sparsity per layer, given the pruning budget.
 Structured pruning, on the other hand, imposes constraints on the sparsification pattern, such as removing rows, columns, or even entire weight tensors.
@@
 (2024); Zhang et al.
-(2024); Dery et al.
-(2024), but without fine-tuning, these methods generally underperform compared to unstructured pruning.
-Our distinction is to introduce a new class of pruning—structured-then-unstructured pruning—and demonstrate its significant advantages for MoEs, surpassing the performance of unstructured-only pruning.
+(2024a); Dery et al.
+(2024), but without fine-tuning, these methods generally underperform when compared to unstructured pruning.
+Our distinction is to introduce a new class of pruning– structured-then-unstructured pruning– and demonstrate its significant advantages for MoEs, surpassing the performance of either method alone.
 This approach differs from previous methods that combine structured and unstructured pruning Kurtic et al.
-(2022), which have not succeeded in outperforming unstructured pruning.
+(2022), which failed to outperform unstructured pruning.
 2.2 Expert Pruning
-Early work on expert pruning was domain-specific, such as in translation MoEs, by keeping most activated experts Kim et al. (2021), or pruning based on gate statistics Koishekenov et al. (2023).
-Lu et al. (2024) introduced a domain-agnostic expert pruning, searching for the best combination of experts to reduce the reconstruction loss, and quantify their criticality in output prediction.
-Our distinction is eliminating the need for expensive combination enumeration, reducing the GPU calls from $O(\frac{k^{n}}{\sqrt{n}})$ to $O(1)$.
+Early work on expert pruning was domain-specific Kim et al.
+(2021); Koishekenov et al.
+(2023); Liu et al.
+(2024), such as in translation MoEs, by keeping most activated experts Kim et al.
+(2021), or pruning based on gate statistics Koishekenov et al.
+(2023).
+Lu et al.
+(2024a) introduced a domain-agnostic expert pruning, searching for the best combination of experts to reduce the reconstruction loss, and quantify their criticality in output prediction.
+Concurrently to our study, Zhang et al.
+(2024b) proposed an efficient expert pruning method.
+Our distinction is two-fold.
+First, we interpolate expert pruning with unstructured pruning to outperform either method alone.
+Second, for scalable expert-level structured pruning, we derive a scalable expert-level structured pruning method with $O(1)$ GPU calls, improving on the $O(\frac{k^{n}}{\sqrt{n}})$ solution enumerating combinatorial pruning.
 2.3 Pruning Robustness
 Robustness in post-hoc pruning is quantified by whether performance is maintained after pruning.
-Training methods aimed at enhancing this robustness include applying group lasso Wen et al.
-(2016); Behnke and Heafield (2021), L1 Han et al.
-(2015); Liu et al.
-(2017), or L0 Louizos et al.
-(2018) regularization to the weights, as well as learning a sparse scaling factor Huang and Wang (2018); Li et al.
-(2020); Gong et al.
-(2022).
-Additionally, Targeted Dropout Gomez et al.
-(2019) demonstrates that applying stochastic dropout to less important parts of the network can increase robustness to pruning.
-Meanwhile, kurtosis of weights Mason-Williams and Dahlqvist (2024) has been used as a proxy of robustness, stating networks with higher weight kurtosis can tolerate higher unstructured pruning ratios.
-Our contribution lies in connecting these findings for designing expert pruning.
-We argue that MoE training inherently enhances robustness to expert-level pruning, and this coarse structured pruning subsequently makes the remaining network more resilient to unstructured pruning.
-In section 5, we argue the former, based on the resemblance of MoE training and targeted dropout, and the latter, inspired by Mason-Williams and Dahlqvist (2024).
+Kurtosis of weights Mason-Williams and Dahlqvist (2024) has been used as a proxy for robustness, with networks showing higher weight kurtosis able to tolerate higher unstructured pruning ratios.
+Our contribution is demonstrating that an expert-pruned network remains robust to additional unstructured pruning, which naturally supports our design of unstructured pruning as the second phase.
 3 Preliminaries: MoE
@@
 $M(x;\mathbf{\theta})=\sum_{i\in\mathcal{T}}r_{i}(x)E(x;\theta_{i})$ (3)
-4 Structured-Then-UNstructured Pruning (STUN)
-4.1 Overview
-Essentially, STUN first performs structured (expert) pruning, until the loss is negligible, then followed by unstructured pruning, such as OWL Yin et al. (2024) or Wanda Sun et al. (2024).
-Our key contribution is to replace combinatorial loss with $O(1)$ expert pruning method, by leveraging latent cluster structure among experts, based on behavioral similarity.
-Specifically, we find clusters of similar experts layer by layer, yielding a total of $\phi nl$ clusters in the whole MoE, where $\phi$ is the pruning ratio, $n$ is the number of experts in each layer, and $l$ is the number of layers in MoE.
+4 Proposed Method
+Figure 2: Overview of our proposed STUN.
+➀ We first remove redundant experts with expert-level structured pruning, then ➁ perform unstructured pruning inside individual experts.
+Black box represents a layer in MoE, and different colors represent different behavioral similarities.
+4.1 Overview: STUN
+Figure 2 overviews our two-phase approach, interpolating structured and unstructured pruning as motivated in Table 1.
+Section 4.2 describes ➀ how we remove redundant experts with expert-level structured pruning with high scalability, then Section 4.3 describes ➁ how we perform unstructured pruning inside individual experts.
+4.2 Expert-level Structured Pruning with $O(1)$ GPU calls
+Now, we describe our expert-level structured pruning with $O(1)$ GPU calls.
+[^3] We focus on GPU cost, since it dominates the CPU cost– For example, the accumulated CPU cost (including the hyperparameter search to meet desired sparsity) required by our algorithm is less than 1 minute, even on 480B Snowflake Arctic.
+Previous solution Lu et al.
+(2024a) minimizes reconstruction loss (Section 4.2.1), requiring GPU call per combination of experts, that is $O(\frac{k^{n}}{\sqrt{n}})$.
+Our key contribution is to approximate this combinatorial reconstruction loss to reduce the number of GPU calls to $O(1)$, by leveraging latent cluster structure among experts, based on behavioral similarity.
+Specifically, we find clusters of similar experts layer by layer, yielding a total of $\phi nl$ clusters in the whole MoE, where $\phi$ is the sparsity, $n$ is the number of experts in each layer, and $l$ is the number of layers in MoE.
 Then we greedily prune every expert but one representative per each cluster.
-Later sections show why our greedy pruning is as effective as its combinatorial counterpart, and how it effectively connects to MoE training beforehand, and unstructured pruning afterwards.
-4.2 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
-Now we will derive our algorithm in section 4.1 formally, starting from the conventional goal of pruning– minimizing the reconstruction loss.
+Later sections show why our greedy pruning is as effective as its combinatorial counterpart.
+4.2.1 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
+We start from the conventional goal of pruning– minimizing the reconstruction loss.
 Reconstruction loss has been employed to assess how closely the pruned model $\theta-\theta_{S}$ without expert set $S$ mirrors the behavior of the unpruned $\theta$  Lu et al.
-(2024).
+(2024a).
 Formally, this loss is quantified by the Frobenius norm of the difference between the original output $M(x;\theta)$ and the output of pruned layer $M(x;\theta-\theta_{S})$, denoted as $\mathcal{E}_{S}$.
@@
 To elaborate, deciding which experts to prune using Eq.
-4 for $|S|=\phi n$ requires (${n\choose{|S|}}\approx O(\frac{k^{n}}{\sqrt{n}})$ forward passes according to Stirling’s approximation, where $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ represents the pruning ratio.
-[^2] In detail, 23951146041928082866135587776380551750 forward passes per layer at minimum, for massive MoEs with $n=128$ experts such as Snowflake Arctic Snowflake (2024).
+4 for $|S|=\phi n$ requires (${n\choose{|S|}}\approx O(\frac{k^{n}}{\sqrt{n}})$ forward passes according to Stirling’s approximation, where $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ represents the sparsity.
 Our distinction is to lower the computation to $O(1)$, without compromising the performance– In fact, as we will elaborate later, we outperform the combinatorial objective.
-4.3 Towards $O(n)$: Probabilistic Interpretation
-As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ as:
-$\textrm{argmax}_{S}P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ (5)
-where $s_{i}$s are the experts included in the expert set $S$, and $P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ is the joint probability of pruning $S$.
-Section 4.2 corresponds to enumerating joint probability from all combinations, requiring ($n\choose{|S|}$ different values, which is compute-intensive.
-When chain rule is applied, Eq.
-5 can be reformulated as follows:
+4.2.2 Towards $O(n)$: Probabilistic Interpretation
+As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ (Eq. 4) as:
 $\textrm{argmax}_{S}\prod_{k}P(X_{k}=s_{k}|X_{1}=s_{1},\\ \cdots,X_{k-1}=s_{k-1})$ (5)
-We propose a greedy optimization for Eq.
-6– We decompose the multiplication of Eq.
-6 and at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability.
-For simplicity, we will omit $X_{k}$s from this point on.
-As our goal is finding the argmax of the probabilities as in Eq.
-6, estimating the rank between them is sufficient, rather than evaluating exact values.
-Such rank estimation can benefit from the latent structure among experts, specifically, a cluster of similar experts in MoE, enabling $P(X_{k}|s_{1},\cdots,s_{k-1})$ calculation without chain-rule multiplications in Eq.
-6.
-Assume we know oracle clusters, $c(E_{i})$, where $c$ is the mapping from an expert to a set of similarly behaving experts identified from the latent clusters.
-When we have knowledge of similar experts, for example, $c(E_{i})=c(E_{j})=\{E_{i},E_{j}\}$ indicating $E_{i}$ and $E_{j}$ are highly similar, we will decide not to prune $E_{i}$ if $E_{j}$ is already pruned.
-That is, if $c(E_{i})\subseteq S_{k}$ then $P(X_{k}=E_{i}|S_{k-1})$ should be lowered by some value $p$, to guide the model against pruning.
-Moreover, $P(E_{i}|S_{k-1})$ should be larger, or rank higher, otherwise.
-To generalize, we will cluster similar experts. Once the cluster of similar experts is finalized, we assign the value $P(E_{i}|S_{k-1})$, as follows:
+Our contribution is greedy optimization without compromise for Eq.
+5– We decompose the multiplication of Eq.
+5 at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability.
+To achieve it, we estimate the rank between the probabilities.
+Such rank estimation can benefit from the latent structure among experts, specifically, a cluster of similar experts in MoE.
+Given cluster mapping $c$ which maps an expert to a set of similarly behaving experts, we assign the value $P(E_{i}|S_{k-1})$, as follows:
 $P(E_{i}|S_{k-1})=\begin{cases}P(E_{i})-p&c(E_{i})\subseteq S_{k}\\ P(E_{i})&\mathrm{otherwise}\end{cases}$ (6)
-We set $p$ as a constant for simplicity. This enables the calculation of all $P(E_{i}|S_{k-1})$ in Eq. 6 from $P(E_{i})$s, which needs only $n$ forwards in total.
+This enables the calculation of all $P(E_{i}|S_{k-1})$ in Eq. 5 from $P(E_{i})$s, which needs only $n$ forwards in total.
 Clustering the Similar Experts
-Our remaining task is to obtain cluster information $c$: One signal is pairwise behavioral similarity $b_{i,j}$, from the pretrained weights $W$ at a minimal cost.
+Our remaining task is to obtain cluster information $c$: One signal is pairwise behavioral similarity $b_{i,j}$, from the pretrained router weights $W$ at a minimal cost.
 Suppose two rows $W_{i}\approx W_{j}$ are similar; then $r_{i}(x)\approx r_{j}(x)$, meaning $E_{i},E_{j}$ tend to be selected by similar inputs, implying similar expertise.
@@
 $b_{i,j}=-\|W_{i}-W_{j}\|_{F}$ (7)
-Next, we generalize pairwise similarity into clusters of experts, such that experts in each cluster $C_{l}$ are highly similar to its representative $\mu_{l}$.
-Formally, the objective of clustering is to minimize the sum of squared errors between $\mu_{l}$ and experts $E_{i}$ in the cluster:
-$\sum_{i\in C_{l}}\sum_{l}\|W_{i}-\mu_{l}\|^{2}$ (9)
-which is an NP-hard problem Megiddo and Supowit (1984); Dasgupta (2008); Aloise et al. (2009).
-Practically, we found that the agglomerative clustering algorithm Sneath and Sokal (1973) performs well.
-[^3] We tried other clustering algorithms in the Appendix.
-Specifically, clusters are initialized as individual experts and then iteratively merged, with a termination condition that prevents the experts within each cluster from being too dissimilar.
-This condition is tuned based on the desired pruning ratio.
-Lastly, if we allow inference on some data, we can improve Eq.
-8 with coactivation statistics $a_{i,j}$, which measure the frequency with which $E_{i},E_{j}$ are selected simultaneously.
-[^4] We normalize $a_{i,j}$ by dividing it with the total coactivations in one layer.
-However, these coactivation statistics depend on the given data, whose distribution may differ from the test data.
-Therefore we combine the two as follows:
-$b_{i,j}=-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$ (10)
-We recap the algorithm in the Appendix (Alg 1).
-4.4 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
-1st-order Taylor Approximation
-While the $O(n)$ approach immensely reduces the cost to obtain the probability distribution, we can further reduce the number of forward passes needed by precalculating $P(E_{i})$s in Eq. 7.
-The key idea is approximating $E_{i}$’s reconstruction loss value $\mathcal{E}_{i}=\|M(x;\theta)-M(x;\theta-\theta_{i})\|_{F}$, and assigning $P(E_{i})$ as some high value $L$ if the reconstruction loss $\mathcal{E}_{i}$ is lowest.
-This neatly estimates the rank between $P(E_{i})$s.
-Though $\mathcal{E}_{i}$ can be approximated via conventional 2nd-order reconstruction methods Hassibi and Stork (1992); Frantar and Alistarh (2023), the size of the hessian matrix increases quadratically with the number of experts, which often yields out-of-memory errors.
-To address this, we propose using a 1st order Taylor approximation.
-To rank the reconstruction loss values, we consider approximating the reconstruction loss when replacing the output from $\theta_{i}$ with some specific expert $\theta_{C}$ in $C=c(E_{i})$ as follows:
-$\mathcal{E}_{i}=\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ (11)
-As the convention of 2nd order Taylor approximation Hassibi and Stork (1992); Frantar and Alistarh (2023), we assume the parameters are near a local minimum.
-Thus, with a small constant $\gamma$, $\|E^{\prime}(\theta_{i})\|<\gamma$, leading to:
-$\sum_{i}\mathcal{E}_{i}<\sum_{i}\gamma\|\theta_{i}-\theta_{C}\|^{2}$ (12)
-whose upper bound in the right-hand side can be minimized when $\theta_{C}=\overline{\theta_{i}}$, where $\bar{~{}}$ denotes the average.
-Therefore, the expert closest to $\bar{\theta_{i}}$ within each cluster has the highest priority to be retained.
-We assign $\mathcal{E}_{i}$ a large number $L>p$ if $E_{i}$ is the closest to $\bar{\theta_{i}}$ from the corresponding cluster $c(E_{i})$, and set it to zero otherwise.
-The same greedy algorithm is applied to optimize Eq.
-6.
-Selective Reconstruction of Experts
-While letting $\theta_{C}$ as the expert closest to $\bar{\theta_{i}}$ successfully minimizes $\sum_{i}\mathcal{E}_{i}$, sometimes we can minimize them further, by replacing the weight of the closest expert $\theta_{C}$ to $\bar{\theta_{i}}$.
-However, blindly doing so is suboptimal, as there is another kind of error to consider.
-The decision boundaries of the next layer are accustomed to the output of $\{E(x;\theta_{i})\}_{i=1}^{|C|}$, but changing the output as $E(x;\theta_{C})=E(x;\bar{\theta_{i}})$ could result in a distribution that the model is unfamiliar with.
-This potential error, which we denote as $\mathcal{E}_{d}$, would be minimized if $\theta_{C}\in\{\theta_{i}\}_{i=1}^{|C|}$.
-To balance these two types of errors, we selectively decide whether to reconstruct.
-We observe that $\sum_{i}\mathcal{E}_{i}$ increases if the total number of clusters in a layer decreases, as this would introduce more $\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ terms.
-Therefore, if the total number of clusters is below a threshold $\kappa$, we use $\theta_{C}=\overline{\theta_{i}}$ to minimize $\sum_{i}\mathcal{E}_{i}$.
-Otherwise, we set $\theta_{C}$ as the expert within the cluster $\{\theta_{i}\}_{i=1}^{|C|}$ closest to the $\overline{\theta_{i}}$, to minimize $\mathcal{E}_{d}$.
-The router weight reconstruction is done similarly, following its corresponding expert.
-The final algorithm is summarized in the Appendix (Alg 2).
-5 Robustness of Structured-Then-Unstructured Pruning
-In this section, we deliver the key insights into why STUN works: MoEs are inherently robust to expert pruning, and the expert-pruned MoE is robust to subsequent post-hoc unstructured pruning.
-MoE Robustness to Structured (Expert) Pruning
-To support this, we find a resemblance between the MoE (Eq.
-3) and the targeted dropout Gomez et al.
-(2019).
-Targeted dropout achieves a higher post-hoc pruning ratio by stochastically dropping out unimportant units or weights during training.
-Formally, in each training step, $M(x;\theta)$ is modified as follows:
-$M(x;\theta-U(\theta)\odot D)$ (13)
-where $U(\theta)$ is a function that extracts only an unimportant part of a given weight $\theta$, and $D$ is a stochastic mask to ensure $U(\theta)$ also has a chance to be updated.
-Gomez et al.
-(2019) designed $U(\theta)$ as the bottom $|\theta|-k$ elements sorted by their absolute values, and $D\sim Bernoulli(\alpha)$ where $\alpha$ is the dropout ratio.
-After training, the performance is better retained when $U(\theta)$ is pruned.
-We recognize the resemblance between the MoE (Eq.
-3) and the targeted dropout (Eq.
-13).
-If we interpret $U(\theta)$ as $\bigcup_{i\notin\mathcal{T}}E(\theta_{i})$, then $M(x;\theta-U(\theta))$ resembles Eq.
-3.
-Moreover, just as $D$ is added in Eq.
-13 to ensure that $U(\theta)$ is updated, MoE ensures that $U(\theta)$ is updated through the stochastic nature of $r(x)$.
-Therefore, after training, MoE is expected to be robust to pruning $U(\theta)$, which, in this context, corresponds to pruning some experts $E(\theta_{i})$s.
-However, identifying $U(\theta)$ is not as straightforward as in Gomez et al.
-(2019).
-In MoE training, $U(\theta)$ is conditioned on each token, such that we need to marginalize token dependency, by identifying unimportant experts across a broad range of inputs.
-This challenge is connected to the problem of finding $S$ in Eq.
-4, which minimizes the reconstruction loss.
-Expert-Pruned MoE Robustness to Unstructured Pruning
-Robustness to unstructured pruning can be estimated by the kurtosis of weights Mason-Williams and Dahlqvist (2024). Kurtosis is expressed as follows:
-$K(\theta)=E\left[\left(\frac{\theta-\mu}{\sigma}\right)^{4}\right]$ (14)
-Suppose the weight of experts $\theta$ follow a zero-meaned Gaussian distribution $\mathcal{N}$.
-Unstructured pruning Sun et al.
-(2024); Yin et al.
-(2024); Das et al.
-(2024); Dong et al.
-(2024), which tends to remove near-zero weights, [^5] The importance score of unstructured pruning typically increases as the absolute value of the weight increases.
-would shift the distribution closer to a bimodal symmetric distribution, whose kurtosis is minimum Darlington (1970).
-As a result, unstructured pruning would lower the kurtosis value, leaving less margin for further unstructured pruning.
-In contrast, coarse structured pruning, such as expert pruning, is less likely to decrease the kurtosis value, since the assumption $\theta\sim\mathcal{N}$ still holds for remaining experts.
-This implies that expert pruning preserves the robustness of unstructured pruning, unlike applying unstructured pruning with a similar pruning ratio.
-6 Experiments
-model sparsity method GSM8K Avg(→) ARC-c ARC-e HellaSwag MMLU
-Arctic 0% unpruned 70.74 68.33 56.91 84.60 66.94 64.86
-40% STUN (w/ OWL) 70.28 67.66 57.68 83.29 64.94 64.75
-OWL 63.76 67.35 56.74 84.13 65.08 63.47
-STUN (w/ Wanda) 69.60 67.64 57.25 83.63 64.86 64.81
-Wanda 64.59 67.54 57.00 84.64 65.19 63.32
-65% STUN (w/ OWL) 43.97 62.67 51.54 80.01 59.91 59.24
-OWL 13.42 56.68 44.37 76.64 53.69 52.02
-Mixtral-8x7B (Instruct) 65% STUN (w/ OWL) 25.09 60.34 48.12 78.79 54.05 60.39
-OWL 1.29 45.20 24.15 49.79 49.27 57.60
-Mixtral-8x22B (Instruct) 70% STUN (w/ OWL) 30.78 60.20 47.95 77.86 55.41 59.56
-OWL 19.64 57.74 45.48 76.60 52.47 56.42
-Table 1: Comparison between STUN and the baselines across various models.
-model sparsity method cost ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
-Mixtral-8x7B (Instruct) 0% unpruned 62.20 87.04 88.50 67.59 68.87 36.60 72.20 76.87 69.98
-25% Ours $O(1)$ 59.30 85.44 88.13 64.42 64.52 35.40 71.84 75.37 68.05
-Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 58.19 84.89 87.34 65.24 62.47 35.60 70.04 75.85 67.45
-Mixtral-8x7B 0% unpruned 57.17 84.01 85.35 64.88 67.88 35.00 70.40 75.93 67.58
-25% Ours $O(1)$ 52.73 81.82 83.09 60.84 63.34 31.60 68.59 72.69 64.34
-Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 51.62 81.94 83.64 61.60 58.72 33.00 67.87 75.37 64.22
-Table 2: Comparing the first component of STUN, the proposed expert pruning, with other baselines.
-(a) Arctic (128x3.66B)
-(b) Mixtral-8x7B
-(c) Mixtral-8x22B
-Figure 2: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various MoEs.
-In this section, we aim to address the following research questions:
-  • RQ1: Does STUN outperform unstructured pruning alone?
-  • RQ2: Does the expert pruning we propose outperform existing baselines?
-  • RQ3: Does STUN favor recent MoE trends with a large number of small experts?
-  • RQ4: Is each component of STUN essential for achieving performance gains?
-  • RQ5: Does STUN generalize to non-MoE models as well?
-6.1 Experimental Settings
-We use Snowflake Arctic Snowflake (2024) as a representative large MoE, with a total of 480B parameters and 128 experts.
-To compare our method with previous works Lu et al.
-(2024), we also experiment with Mixtral Jiang et al.
-(2024).
-Tasks and Datasets
-In contrast to previous unstructured pruning studies Sun et al.
-(2024); Yin et al.
-(2024), we also evaluate on the NLG task, GSM8K Cobbe et al.
-(2021), where maintaining performance proves to be much more challenging (see Table 1).
-We further assess performance on four NLU tasks– ARC-challenge and ARC-easy Clark et al.
-(2018), HellaSwag Zellers et al.
-(2019), and MMLU Hendrycks et al.
-(2021).
-When comparing with expert pruning methods, following previous work Lu et al.
-(2024), we also conduct a zero-shot evaluation on BoolQ Wang et al.
-(2019), OpenBookQA Mihaylov et al.
-(2018), RTE Wang et al.
-(2018), WinoGrande Sakaguchi et al.
-(2021).
-To provide some data for inference, we employ the C4 dataset Raffel et al. (2020), following the baselines Yin et al. (2024); Sun et al. (2024); Lu et al. (2024).
-Implementation Details
-We explore the values of $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for Snowflake Arctic, the largest MoE in our experiments, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, meaning no GPU calls are required for expert pruning.
-We evaluate methods using the lm-evaluation-harness Gao et al.
-(2021).
-Due to the model size, we use 4-bit quantization Dettmers et al.
-(2023) for experiments with Mixtral-8x22B and Arctic.
-The expert pruning ratios are set to 20%, 12.5%, and 10% for Snowflake Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively.
-We use $\kappa=3$ for selective reconstruction.
-More detailed implementation information is provided in the Appendix.
-6.2 Experimental Results
-6.2.1 RQ1: STUN Outperforms Unstructured Pruning
-Table 1 describes that our proposed (STUN) significantly outperforms the unstructured pruning methods. We emphasize that we use the same unstructured pruning approach for all for fair comparison.
-For example, with 40% of sparsity for the Arctic, STUN neatly retains the original GSM8K performance, while unstructured pruning results in a noticeable performance drop.
-This is consistent for different unstructured pruning methods, Wanda, as well.
-For 65% of sparsity for Mixtral-8x7B-Instruct, STUN’s GSM8K performance is nearly 20 times better than that of unstructured pruning.
-In the ARC-challenge, the unstructured pruning performance falls below the random-guess accuracy of 25.02 Clark et al.
-(2018), whereas STUN maintains a significantly higher performance, achieving twice the score.
-6.2.2 RQ2: Our $O(1)$ Expert Pruning Outperforms Existing $O(\frac{k^{n}}{\sqrt{n}})$
-Table 2 shows that our proposed $O(1)$ expert pruning method is highly effective, outperforming the previous $O(\frac{k^{n}}{\sqrt{n}})$ solution.
-This is because we derive the latent structure from the pretrained MoE, while the previous work Lu et al.
-(2024) relies solely on the given calibration data.
-This validates our design of $O(1)$ in section 4.
-6.2.3 RQ3: STUN Favors Large Number of Small Experts
-Figure 2 illustrates the trend of STUN in 3 different MoEs.
-The performance gap between STUN and unstructured pruning increases as the MoE has more experts with small sizes (from (c) to (a)).
-This is because having more experts, rather than having fewer but larger ones, provides greater flexibility to our expert pruning.
-Notably, MoEs with a large number of small experts are favored in recent models He (2024).
-Cluster (§$\S$ 4.3) Reconstruct (§$\S$ 4.4) LM-eval Avg
-Ours Ours 59.58
-DSatur Ours 58.59
-Ours Always 57.60
-Ours Never 59.22
-Table 3: Ablation experiments for the first component of STUN, the proposed expert pruning.
-6.2.4 RQ4: Ablation
-To validate our design of expert pruning in sections 4.3 and 4.4, we evaluate alternative approaches to expert-prune Mixtral-8x7B at 50% sparsity.
-Table 3 confirms that our design choices are valid.
-Our agglomerative clustering algorithm outperforms the DSatur algorithm, an alternative clustering algorithm we discuss in the Appendix.
-Additionally, selective reconstruction proves superior to always or never reconstructing, as shown in the last two rows.
-Detailed evaluation results are provided in the Appendix.
-6.2.5 RQ5: STUN Outperforms Unstructured Pruning in non-MoEs
-To investigate whether STUN is generalizable to non-MoE as well, we employ a state-of-the-art structured pruning algorithm for non-MoE model, namely, LLM-surgeon van der Ouderaa et al.
-(2024) with 5% sparsity before performing unstructured pruning, which is OWL in our case.
-Figure 3 illustrates that such STUN outperforms unstructured pruning.
-(a) Llama-2 7B
-(b) Llama-2 13B
-Figure 3: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various non-MoEs.
-7 Conclusion
-In this paper, we proposed STUN for MoEs, which outperforms unstructured-only pruning.
-Although this may seem counterintuitive, we provide both theoretical and empirical evidence demonstrating why designing expert pruning before unstructured pruning is beneficial.
-Since existing expert pruning requires $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, we proposed $O(1)$ complexity, which outperforms the more expensive approaches.
-Our experimental results show STUN  is highly effective– For 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to.
-Limitation
-Since our method utilizes unstructured pruning in the second stage, we share the same disadvantages with unstructured pruning, that is, it may need specialized hardware for acceleration.
-However, it is shown that general-purpose hardware, such as CPU, can successfully accelerate unstructure-pruned networks NeuralMagic (2021).
-Therefore we have the same potential as existing unstructured pruning methods– with the support from the community, our method will scale up to existing accelerators.
-Appendix
-Algorithm of $O(1)$ Expert Pruning
-Algorithms 1 and 2 describe the expert pruning we proposed. Note that we do not introduce any GPU inference in both two algorithms.
+which can be improved with coactivation statistics $a_{i,j}$, if we allow some inference cost.
+As a result, we illustrate our clustering algorithm in Alg 1, whose detailed derivation can be found in Appendix A.1.
 Algorithm 1 Expert Clustering Algorithm
@@
 $\lambda_{1},\lambda_{2}\leftarrow$ Hyperparameter for behavioral similarity
-$t\leftarrow$ Threshold to determine pruning ratio
+$t\leftarrow$ Threshold to determine sparsity
+$c\leftarrow$ The mapping from expert to cluster of the similar experts
 […]
@@
 return $c$
+4.2.3 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
+While the previous section immensely reduces the cost to obtain the probability distribution to $O(n)$ by requiring only $P(E_{i})$s, we can further reduce the number of forward passes– We aim to remove the GPU calls for $P(E_{i})$, which is needed as in Eq.
+10.
+The key idea is approximating $E_{i}$’s reconstruction loss value $\mathcal{E}_{i}=\|M(x;\theta)-M(x;\theta-\theta_{i})\|_{F}$.
+To address this, with 1st order Taylor approximation, we find the expert closest to $\bar{\theta_{i}}$ within each cluster has the highest priority to be retained.
+We assign ranks similarly, and the same greedy algorithm is applied to optimize Eq.
+5.
+Additionally, we selectively reconstruct the expert.
+The final algorithm is summarized in Alg 2, which is described in detail in Appendix A.2.
 Algorithm 2 Our $O(1)$ Expert Pruning
@@
 return $r$
+4.3 Unstructured Pruning on Expert-pruned Model
+Our main conjecture for STUN is intra-expert sparsity yet remains intact after the first phase.
+Specifically, we propose to pursue fine-grained sparsity within the remaining experts, by leveraging unstructured pruning methods designed for general LLM , such as OWL Yin et al.
+(2024) or Wanda Sun et al.
+(2024).
+Now we theoretically verify our conjecture: intra-expert sparsity remains high, or, even higher, after the expert pruning phase.
+In other words, we explain why performing unstructured pruning after expert pruning is better than continuously performing unstructured pruning only.
+Intra-expert sparsity, formally speaking, robustness to unstructured pruning, can be estimated by the kurtosis of weights Mason-Williams and Dahlqvist (2024). Kurtosis is expressed as follows:
+$K(\theta)=E\left[\left(\frac{\theta-\mu}{\sigma}\right)^{4}\right]$ (8)
+Suppose the weight of experts $\theta$ follow a zero-meaned Gaussian distribution $\mathcal{N}$.
+Unstructured pruning Sun et al.
+(2024); Yin et al.
+(2024); Das et al.
+(2024); Dong et al.
+(2024), which tends to remove near-zero weights, [^4] The importance score of unstructured pruning typically increases as the absolute value of the weight increases.
+would shift the distribution closer to a bimodal symmetric distribution, whose kurtosis is minimum Darlington (1970).
+As a result, unstructured pruning would lower the kurtosis value, leaving less margin for further unstructured pruning.
+In contrast, coarse structured pruning, such as expert pruning, is less likely to decrease the kurtosis value, since the assumption $\theta\sim\mathcal{N}$ still holds for remaining experts.
+This implies that expert pruning preserves the robustness of unstructured pruning, unlike applying unstructured pruning with a similar sparsity.
+[^5] In our experiments, the kurtosis increased from 14248 to 15623 after expert pruning.
+5 Experiments
+model sparsity method GSM8K Avg(→) ARC-c ARC-e HellaSwag MMLU
+Arctic 0% unpruned 70.74 68.33 56.91 84.60 66.94 64.86
+40% STUN (w/ OWL) 70.28 67.66 57.68 83.29 64.94 64.75
+OWL 63.76 67.35 56.74 84.13 65.08 63.47
+STUN (w/ Wanda) 69.60 67.64 57.25 83.63 64.86 64.81
+Wanda 64.59 67.54 57.00 84.64 65.19 63.32
+65% STUN (w/ OWL) 43.97 62.67 51.54 80.01 59.91 59.24
+OWL 13.42 56.68 44.37 76.64 53.69 52.02
+Mixtral-8x7B (Instruct) 65% STUN (w/ OWL) 25.09 60.34 48.12 78.79 54.05 60.39
+OWL 1.29 45.20 24.15 49.79 49.27 57.60
+LLM-Pruner 1.29 31.74 22.78 45.96 35.09 23.13
+Mixtral-8x22B (Instruct) 70% STUN (w/ OWL) 30.78 60.20 47.95 77.86 55.41 59.56
+OWL 19.64 57.74 45.48 76.60 52.47 56.42
+Table 2: Comparison between STUN and the baselines across various models.
+model sparsity method cost Avg
+Mixtral-8x7B (Instruct) 0% unpruned 69.98
+25% Ours $O(1)$ 68.05
+Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 67.45
+Mixtral-8x7B 0% unpruned 67.58
+25% Ours $O(1)$ 64.34
+Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 64.22
+Table 3: Comparing the average performance of 8 tasks of the proposed expert pruning, with other baselines.
+sparsity ARC-C BoolQ HellaSwag MMLU RTE WinoGrande Avg
+Unpruned 0% 59.4 84.2 84.0 67.9 70.4 75.6 71.5
+Ours 25% 55.6 83.1 81.1 63.3 68.6 72.7 70.7
+SEER-MoE Muzio et al. (2024) 25% - - - 56.7 - - -
+Expert Drop He et al. (2024) 25% 53.2 77.7 80.5 52.2 55.6 76.8 66.0
+Layer Drop He et al. (2024) 25% 47.7 85.3 75.2 67.3 69.7 74.6 70.0
+Table 4: Comparison of efficient expert pruning methods on Mixtral-8x7B.
+sparsity cost GSM8K
+unpruned 0% 63.46
+Ours 25% $O(1)$ 53.22
+Lu et al. (2024a) 25% $O(\frac{k^{n}}{\sqrt{n}})$ 48.52
+Table 5: GSM8K accuracy comparison with baseline on Mixtral-8x7B-Instruct.
+5.1 Experimental Settings
+We use Snowflake Arctic Snowflake (2024) as a representative large MoE, with a total of 480B parameters and 128 experts.
+To compare our method with previous works Lu et al.
+(2024a), we also experiment with Mixtral Jiang et al.
+(2024).
+Tasks and Datasets
+In contrast to previous unstructured pruning studies Sun et al.
+(2024); Yin et al.
+(2024), we also evaluate the NLG task, GSM8K Cobbe et al.
+(2021), where maintaining performance proves to be much more challenging (see Table 2; Appendix F).
+We further assess performance on four NLU tasks– ARC-challenge and ARC-easy Clark et al.
+(2018), HellaSwag Zellers et al.
+(2019), and MMLU Hendrycks et al.
+(2021).
+When comparing with expert pruning methods, following previous work Lu et al.
+(2024a), we also conduct a zero-shot evaluation on BoolQ Wang et al.
+(2019), OpenBookQA Mihaylov et al.
+(2018), RTE Wang et al.
+(2018), WinoGrande Sakaguchi et al.
+(2021).
+To provide some data for inference, we employ the C4 dataset Raffel et al. (2020), following the baselines Yin et al. (2024); Sun et al. (2024); Lu et al. (2024a).
 Implementation Details
+We explore the values of $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for Snowflake Arctic, the largest MoE in our experiments, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, meaning no GPU calls are required for expert pruning.
+The sparsity for expert-level structured pruning is set to 20%, 12.5%, and 10% for Snowflake Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively.
+We use $\kappa=3$ for selective reconstruction.
+More detailed implementation information and hyperparameter decisions are provided in Appendix.
+5.2 Experimental Results
+5.2.1 RQ1: STUN Outperforms Unstructured Pruning
+Table 2 describes that our proposed (STUN) significantly outperforms the unstructured pruning methods. We emphasize that we use the same unstructured pruning approach for all for fair comparison.
+For example, with 40% of sparsity for the Arctic, STUN neatly retains the original GSM8K performance, while unstructured pruning results in a noticeable performance drop.
+This is consistent for different unstructured pruning methods, Wanda, as well.
+As the sparsity increases, the table shows that STUN can maintain the original performance much better than the baselines– For 65% sparsity, STUN’s GSM8K performance is nearly 30%p better than that of unstructured pruning.
+With different models, we observe different gaps but a similar trend– For 65% of sparsity for Mixtral-8x7B-Instruct, STUN’s GSM8K performance is nearly 20 times better than that of unstructured pruning.
+In the ARC-challenge, the unstructured pruning performance falls below the random-guess accuracy of 25.02 Clark et al.
+(2018), whereas STUN maintains a significantly higher performance, achieving twice the score.
+Mixtral-8x22B shows a similar trend.
+sparsity $O(n)$ $O(1)$
+25% 63.97 64.34
+50% 59.90 59.58
+Table 6: Comparing the average performance of 8 tasks of our $O(n)$ and $O(1)$ algorithms for expert pruning on Mixtral-8x7B.
+(a) Arctic (128x3.66B)
+(b) Mixtral-8x7B
+(c) Mixtral-8x22B
+(d) OLMoE-7B
+Figure 3: Comparing STUN and unstructured pruning for various MoEs.
+(e) Llama-2 7B
+(f) Llama-2 13B
+Figure 4: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various non-MoEs.
+5.2.2 RQ2: Our $O(1)$ Expert Pruning Outperforms Existing Methods
+Tables 3 and 5 show that our proposed $O(1)$ expert pruning method is highly effective, outperforming the previous $O(\frac{k^{n}}{\sqrt{n}})$ solution.
+This is because we derive the latent structure from the pretrained MoE, while the previous work Lu et al.
+(2024a) relies solely on the given calibration data.
+This validates our design of $O(1)$ in section 4.
+The detailed results are in Appendix E.
+Table 4 shows that our expert pruning outperforms other efficient pruning methods.
+Compared to SEER-MoE, our pruning clearly outperforms the performance in MMLU, by a substantial margin.
+Moreover, ours achieves a higher average performance compared to the results reported by He et al.
+(2024).
+[^6] Note that SEER-MoE Muzio et al.
+(2024) only reveals the performance of MMLU, and we used different metrics following He et al.
+(2024).
+Moreover, Table 6 shows that our $O(1)$ expert pruning method achieves similar performance to our $O(n)$ method. This supports our choice to use the $O(1)$ method, which is more efficient.
+Train GPU cost GPUs #
+Lu et al. (2024a) 0 infeasible [^7] In detail, 23951146041928082866135587776380551750 forward passes per layer at minimum. > 8
+STUN (w/ OWL) 0 1.12h 1
+OWL Yin et al. (2024) 0 1.12h 1
+STUN (w/ Wanda) 0 0.58h 1
+Wanda Sun et al. (2024) 0 0.58h 1
+Table 7: Cost comparison of diverse pruning methods with Snowflake-Arctic. Train: training cost, GPU #: number of GPUs required for pruning.
+5.2.3 RQ3: STUN Adapts Large Number of Small Experts
+Figure 3 illustrates the trend of STUN in different MoEs.
+The performance gap between STUN and unstructured pruning increases as the MoE has more experts with small sizes (from (c) to (a)).
+This is because having more experts, rather than having fewer but larger ones, provides greater flexibility to our expert pruning.
+Notably, MoEs with a large number of small experts are favored in recent models He (2024).
+5.2.4 RQ4: STUN Outperforms Unstructured Pruning in non-MoEs
+To investigate whether STUN is generalizable to non-MoE as well, we employ a state-of-the-art structured pruning algorithm for non-MoE models, namely, LLM-surgeon van der Ouderaa et al.
+(2024) with 5% sparsity before performing unstructured pruning, which is OWL in our case.
+Figure 4 illustrates that such STUN outperforms unstructured pruning.
+6 Cost Analysis
+Table 7 shows the cost comparison between diverse pruning methods.
+While none of the pruning methods require training cost, Lu et al.
+(2024a) is infeasible due to its prohibitive number of forward passes in GPUs.
+Due to the efficiency of proposed expert pruning, STUN is as efficient as the unstructured pruning method it uses, making it a feasible pruning method even for large MoEs, such as Snowflake-Arctic.
+7 Conclusion
+In this paper, we proposed STUN– an interpolation between structured and unstructured pruning, leveraging both inter- and intra-expert sparsity.We provide both theoretical and empirical evidence demonstrating why designing expert pruning before unstructured pruning is beneficial.
+Limitation
+Since our method utilizes unstructured pruning in the second stage, we share the same disadvantages with unstructured pruning, that is, on some hardware, the acceleration may not be trivial.
+However, it is shown that some hardware, such as CPU, can successfully accelerate unstructure-pruned networks NeuralMagic (2021), and ongoing research is actively developing methods to achieve similar speedups on GPUs Mishra et al.
+(2021); Zhao et al.
+(2024).
+With a substantial body of work dedicated to unstructured pruning Frantar and Alistarh (2023); Das et al.
+(2024); Dong et al.
+(2024); Sun et al.
+(2024); Yin et al.
+(2024); Li et al.
+(2024); Hu et al.
+(2024); Lu et al.
+(2024b), we believe this approach remains highly relevant and leads to practical performance improvements with ongoing hardware support.
+We already have shown our method works in versatile extreme settings, such as non-MoE models, or high sparsity, etc.
+We leave more extensive evaluations, such as performance under highly imbalanced or skewed data distributions, as future work.
+Appendix A Derivation From $O(\frac{k^{n}}{\sqrt{n}})$ to $O(1)$
+A.1 Towards $O(n)$: Probabilistic Interpretation
+As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ (Eq. 4) as:
+$\textrm{argmax}_{S}P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ (9)
+where $s_{i}$s are the experts included in the expert set $S$, and $P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ is the joint probability of pruning $S$.
+One intuitive way to design $P$ so that Eq.
+9 yields the same $S$ to minimize Eq.
+4 is as follows:
+$P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})=\frac{1}{Z}\cdot\frac{1}{\mathcal{E}_{S}}$ (10)
+where $Z$ is the normalization factor. Each joint probability needs $O(1)$ GPU calls, since $\mathcal{E}_{S}$ needs the output of $M(x;\theta-\theta_{S})$.
+Section 4.2.1 corresponds to enumerating joint probability from all combinations, requiring ($n\choose{|S|}$ different values, which is compute-intensive.
+When chain rule is applied, Eq.
+9 can be reformulated as follows:
+$\textrm{argmax}_{S}\prod_{k}P(X_{k}=s_{k}|X_{1}=s_{1},\\ \cdots,X_{k-1}=s_{k-1})$ (11)
+Our contribution is greedy optimization without compromise for Eq.
+11– We decompose the multiplication of Eq.
+11 at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability.
+For simplicity, we will omit $X_{k}$s from this point on.
+As our goal is finding the argmax of the probabilities as in Eq.
+11, estimating the rank between them is sufficient, rather than evaluating exact values.
+Such rank estimation can benefit from the latent structure among experts, specifically, a cluster of similar experts in MoE, enabling $P(X_{k}|s_{1},\cdots,s_{k-1})$ calculation without chain-rule multiplications in Eq.
+11.
+Assume we know oracle clusters, $c(E_{i})$, where $c$ is the mapping from an expert to a set of similarly behaving experts identified from the latent clusters.
+When we have knowledge of similar experts, for example, $c(E_{i})=c(E_{j})=\{E_{i},E_{j}\}$ indicating $E_{i}$ and $E_{j}$ are highly similar, we will decide not to prune $E_{i}$ if $E_{j}$ is already pruned.
+That is, if $c(E_{i})\subseteq S_{k}$ then $P(X_{k}=E_{i}|S_{k-1})$ should be lowered by some value $p$, to guide the model against pruning.
+Moreover, $P(E_{i}|S_{k-1})$ should be larger, or rank higher, otherwise.
+To generalize, we will cluster similar experts. Once the cluster of similar experts is finalized, we assign the value $P(E_{i}|S_{k-1})$, as follows:
+$P(E_{i}|S_{k-1})=\begin{cases}P(E_{i})-p&c(E_{i})\subseteq S_{k}\\ P(E_{i})&\mathrm{otherwise}\end{cases}$ (12)
+We set $p$ as a constant for simplicity. This enables the calculation of all $P(E_{i}|S_{k-1})$ in Eq. 11 from $P(E_{i})$s, which needs only $n$ forwards in total.
+Clustering the Similar Experts
+Our remaining task is to obtain cluster information $c$: One signal is pairwise behavioral similarity $b_{i,j}$, from the pretrained weights $W$ at a minimal cost.
+Suppose two rows $W_{i}\approx W_{j}$ are similar; then $r_{i}(x)\approx r_{j}(x)$, meaning $E_{i},E_{j}$ tend to be selected by similar inputs, implying similar expertise.
+Thus, the behavioral similarity $b_{i,j}$ between two experts $E_{i},E_{j}$ can be obtained as follows:
+$b_{i,j}=-\|W_{i}-W_{j}\|_{F}$ (13)
+Next, we generalize pairwise similarity into clusters of experts, such that experts in each cluster $C_{l}$ are highly similar to its representative $\mu_{l}$.
+Formally, the objective of clustering is to minimize the sum of squared errors between $\mu_{l}$ and experts $E_{i}$ in the cluster:
+$\sum_{i\in C_{l}}\sum_{l}\|W_{i}-\mu_{l}\|^{2}$ (14)
+which is an NP-hard problem Megiddo and Supowit (1984); Dasgupta (2008); Aloise et al. (2009).
+Practically, we found that the agglomerative clustering algorithm Sneath and Sokal (1973) performs well.
+[^8] We tried other clustering algorithms in the Appendix.
+Specifically, clusters are initialized as individual experts and then iteratively merged, with a termination condition that prevents the experts within each cluster from being too dissimilar.
+This condition is tuned based on the desired sparsity.
+Lastly, if we allow inference on some data, we can improve Eq.
+13 with coactivation statistics $a_{i,j}$, which measure the frequency with which $E_{i},E_{j}$ are selected simultaneously.
+[^9] We normalize $a_{i,j}$ by dividing it with the total coactivations in one layer.
+However, these coactivation statistics depend on the given data, whose distribution may differ from the test data.
+Therefore, we combine the two as follows:
+$b_{i,j}=-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$ (15)
+We recap the algorithm in the Appendix (Alg 1).
+A.2 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
+1st-order Taylor Approximation
+While previous section immensely reduces the cost to obtain the probability distribution to $O(n)$ by requiring only $P(E_{i})$s, we can further reduce the number of forward passes– We aim to remove the GPU calls for $P(E_{i})$, which is needed as in Eq.
+10.
+The key idea is approximating $E_{i}$’s reconstruction loss value $\mathcal{E}_{i}=\|M(x;\theta)-M(x;\theta-\theta_{i})\|_{F}$, and assigning $P(E_{i})$ as some high value $L$ if the reconstruction loss $\mathcal{E}_{i}$ is lowest.
+This neatly estimates the rank between $P(E_{i})$s.
+Though $\mathcal{E}_{i}$ can be approximated via conventional 2nd-order reconstruction methods Hassibi and Stork (1992); Frantar and Alistarh (2023), the size of the hessian matrix increases quadratically with the number of experts, which often yields out-of-memory errors.
+To address this, we propose using a 1st order Taylor approximation.
+To rank the reconstruction loss values, we consider approximating the reconstruction loss when replacing the output from $\theta_{i}$ with some specific expert $\theta_{C}$ in $C=c(E_{i})$ as follows:
+$\mathcal{E}_{i}=\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ (16)
+As the convention of 2nd order Taylor approximation Hassibi and Stork (1992); Frantar and Alistarh (2023), we assume the parameters are near a local minimum.
+Thus, with a small constant $\gamma$, $\|E^{\prime}(\theta_{i})\|<\gamma$, leading to:
+$\sum_{i}\mathcal{E}_{i}<\sum_{i}\gamma\|\theta_{i}-\theta_{C}\|^{2}$ (17)
+whose upper bound in the right-hand side can be minimized when $\theta_{C}=\overline{\theta_{i}}$, where $\bar{~{}}$ denotes the average.
+Therefore, the expert closest to $\bar{\theta_{i}}$ within each cluster has the highest priority to be retained.
+We assign $\mathcal{E}_{i}$ a large number $L>p$ if $E_{i}$ is the closest to $\bar{\theta_{i}}$ from the corresponding cluster $c(E_{i})$, and set it to zero otherwise.
+The same greedy algorithm is applied to optimize Eq.
+11.
+Selective Reconstruction of Experts
+While letting $\theta_{C}$ as the expert closest to $\bar{\theta_{i}}$ successfully minimizes $\sum_{i}\mathcal{E}_{i}$, sometimes we can minimize them further, by replacing the weight of the closest expert $\theta_{C}$ to $\bar{\theta_{i}}$.
+However, blindly doing so is suboptimal, as there is another kind of error to consider.
+The decision boundaries of the next layer are accustomed to the output of $\{E(x;\theta_{i})\}_{i=1}^{|C|}$, but changing the output as $E(x;\theta_{C})=E(x;\bar{\theta_{i}})$ could result in a distribution that the model is unfamiliar with.
+This potential error, which we denote as $\mathcal{E}_{d}$, would be minimized if $\theta_{C}\in\{\theta_{i}\}_{i=1}^{|C|}$.
+To balance these two types of errors, we selectively decide whether to reconstruct.
+We observe that $\sum_{i}\mathcal{E}_{i}$ increases if the total number of clusters in a layer decreases, as this would introduce more $\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ terms.
+Therefore, if the total number of clusters is below a threshold $\kappa$, we use $\theta_{C}=\overline{\theta_{i}}$ to minimize $\sum_{i}\mathcal{E}_{i}$.
+Otherwise, we set $\theta_{C}$ as the expert within the cluster $\{\theta_{i}\}_{i=1}^{|C|}$ closest to the $\overline{\theta_{i}}$, to minimize $\mathcal{E}_{d}$.
+The router weight reconstruction is done similarly, following its corresponding expert.
+Appendix B Implementation Details
 We probe $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for the Snowflake Arctic, which is the biggest MoE we deal with, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, which means no GPU calls is needed for expert pruning.
@@
 (2023) for experiments with Mixtral-8x22B and Arctic, due to their model size.
-We use 20%, 12.5%, and 10% for Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively, as the expert pruning ratio for STUN.
+We use 20%, 12.5%, and 10% for Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively, as the expert sparsity for STUN.
 These are the maximum values among 10, 12.5, 20, 25, and 35%, with minimum performance loss.
@@
 Evaluation is done only once, since we introduce no randomness in our experiment.
-Other Clustering Algorithms
+Appendix C Other Clustering Algorithms
 We also considered DSatur Brélaz (1979) as a clustering algorithm for Eq.
@@
 $e_{i,j}=\begin{cases}1&b_{i,j}>=t_{DSatur}\\ \infty&\mathrm{otherwise}\end{cases}$ (18)
-where $t_{DSatur}$ is some threshold to control the pruning ratio of MoE.
-Ablation Studies Results
-Detailed per-task performance of ablation studies are described in Tables 4, 5.
+where $t_{DSatur}$ is some threshold to control the sparsity of MoE.
+Cluster Reconstruct LM-eval Avg
+Ours Ours 59.58
+DSatur Ours 58.59
+Ours Always 57.60
+Ours Never 59.22
+Table 8: Ablation experiments for the first component of STUN, the proposed expert pruning.
+Appendix D Ablation Studies
+To validate our design of expert pruning in sections A.1 and A.2, we evaluate alternative approaches to expert-prune Mixtral-8x7B at 50% sparsity.
+Table 8 confirms that our design choices are valid.
+Our agglomerative clustering algorithm outperforms the DSatur algorithm, an alternative clustering algorithm we discuss in the Appendix.
+Additionally, selective reconstruction proves superior to always or never reconstructing, as shown in the last two rows.
+Detailed per-task performance of ablation studies are described in Tables 9, 10.
+Detailed hyperparameter ablation studies are described in Tables 11, 12.
+Results with applying expert-pruning only are provided in Table 14.
 ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
@@
 Table 10: Selective reconstruction outperforms the baselines.
+$\lambda_{1}$ $\lambda_{2}$ GSM8K
+Mixtral-8x7B 1 0 63.38
+0 1 58.53
+1 1 60.42
+Mixtral-8x22B 1 0 81.50
+0 1 81.58
+1 1 80.52
+Table 11: Comparison of different $\lambda_{1}$, $\lambda_{2}$ configurations. For the Arctic, we only consider the cheapest, ($\lambda_{1}$, $\lambda_{2}$) = (1,0).
+sparsity GSM8K
+Arctic 10% 70.74
+20% 69.90
+35% 60.05
+Mixtral-8x7B 12.50% 63.38
+25% 53.22
+Mixtral-22B 10% 81.58
+12.50% 79.91
+25% 74.07
+Table 12: Comparison of different sparsity of our proposed expert-level pruning. ($\lambda_{1}$, $\lambda_{2}$) is probed among (1,0),(0,1),(1,1), except for 10% and 20% of Arctic.
+model sparsity method cost ARC-c ARC-e BoolQ HellaSwag MMLU OBQA RTE WinoGrande Avg
+Mixtral-8x7B (Instruct) 0% unpruned 62.20 87.04 88.50 67.59 68.87 36.60 72.20 76.87 69.98
+25% Ours $O(1)$ 59.30 85.44 88.13 64.42 64.52 35.40 71.84 75.37 68.05
+Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 58.19 84.89 87.34 65.24 62.47 35.60 70.04 75.85 67.45
+Mixtral-8x7B 0% unpruned 57.17 84.01 85.35 64.88 67.88 35.00 70.40 75.93 67.58
+25% Ours $O(1)$ 52.73 81.82 83.09 60.84 63.34 31.60 68.59 72.69 64.34
+Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 51.62 81.94 83.64 61.60 58.72 33.00 67.87 75.37 64.22
+Table 13: Comparing the first component of STUN, the proposed expert pruning, with other baselines.
+sparsity GSM8K Avg (→) ARC-c ARC-e HellaSwag MMLU
+Arctic 20% 69.90 67.95 57.00 83.71 65.70 65.39
+Mixtral-8x7B 12.50% 63.38 70.65 61.18 86.66 66.80 67.97
+Mixtral-8x22B 10% 81.58 71.78 61.18 85.82 66.62 73.52
+Table 14: Performance when applying only our expert pruning.
+Appendix E Detailed Results for RQ2
+Table 13 describes the per-task performance of RQ2.
+Appendix F Why GSM8K is Harder
+The nature of the GSM8K task, which is a generation task, accounts for this discrepancy.
+A random generation baseline on GSM8K would achieve 0% accuracy, making it far more challenging to maintain performance.
+In contrast, ARC, HellaSwag, and MMLU are multiple-choice tasks where random baselines can achieve reasonable accuracy by comparing the perplexity of different completion options.
+As a result, maintaining performance on GSM8K is considerably harder, a challenge often overlooked in previous works.
+This explains the more pronounced difference in performance between STUN and the baselines on GSM8K.